OpenStudy (anonymous):

Find the period of the function. y = -3 cos x

OpenStudy (anonymous):

the period would be 2pi since there is no b term, correct?

OpenStudy (debbieg):

Right. Period of y=Acos(Bx) = \(2\pi/B\). Here B=1, so per=\(2\pi\).

OpenStudy (anonymous):

thank you!

OpenStudy (debbieg):

yw